Circular No. 7456

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cv IN PGC 39222 W. D. Li, University of California at Berkeley, reports the discovery by

M. Schwartz, Cottage Grove, OR, of an apparent supernova (mag ∼ 16.1) on unfiltered CCD images obtained in the course of the Tenagra Observatory Supernova Search with the 0.35-m Tenagra-I automated telescope on July 11.2 and 12.2 UT. SN 2000cv is located at α = 12^{h}15^{m}42^{s}.04, δ = +61^{o}53′24′′.6 (equinox 2000.0). SN 2000cv was also confirmed in an unfiltered CCD image obtained by Li with the Katzman Automatic Imaging Telescope (KAIT) on July 12.2 (mag ∼ 16.0), from which Li measured the position end figures as 42^{s}.21, 24′′.0, which is 9′′.6 east and 1′′.2 north of the nucleus of PGC 39222. Digital Sky Survey images of the field from 1955 (limiting mag ∼ 20.0), 1993 (limiting mag 21.5), and 1994 (limiting mag ∼ 21.0) show no stellar object at the position of SN 2000cv.

SUPERNOVA 2000cw IN MCG +5-56-007 M. Modjaz, A. B. Aazami, and W. D. Li, University of California at

Berkeley, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the KAIT of an apparent supernova on unfiltered images taken on July 14.5 (mag ∼ 17.4) and 15.5 UT (mag ∼ 17.2). The new object is located at α = 23^{h}47^{m}25^{s}.10, δ = +28^{o}23′15′′.8 (equinox 2000.0), which is 8′′.4 east and 21′′.2 south of the nucleus of MCG +5-56-007. A KAIT image of the same field taken on July 5.5 showed nothing at the position of SN 2000cw (limiting mag ∼ 19.0).

XTE J1859+226 J. A. Tomsick and W. A. Heindl, Center for Astrophysics and Space

Sciences, University of California at San Diego, write: “An RXTE observation of the x-ray transient XTE J1859+226 made on July 8.5 UT indicates a 2.5–20-keV flux level of 3 × 10^{−11} erg cm^{−2} s^{−1}, which is well above the expected quiescent x-ray flux level. The higher-than-expected x-ray flux may be related to the optical minioutburst reported on IAUC 7451. The energy spectrum is well-described by a power law with a photon index of 2.03 ± 0.14. There is evidence for two short (< 16 s) x-ray dips during the 3000-s RXTE observation, possibly indicating a moderately high binary inclination. During the deepest dip, the flux drops to 33 ± 19 percent of the mean flux level in the energy band 2.9–10.5 keV. Further x-ray observations are required to determine if the dips are related to the binary orbit.”

2000 July 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green